In a study of 32 students grade points averages (GPA) the following data were obtained.construct a frequency distribution using seven classes.(note:A=4, B=03, C=2, D=1, F=0.) (The information in the exercise will be used again for exercise 2-22 and 3-22).

(3.2,1.1,0.6,2.8,2.6,3.0). (2.0,3.5,4.0,2.6,2.3,1.7). (3.3,1.9,3.5,1.6,3.8,4.0). (2.7,1.7,2.3,1.6,2.1,1.2). (2.1,0.8,1.6,2.4,2.9,3.1).
(3.9,2.6).
1
Expert's answer
2020-01-20T09:41:25-0500

Minimum value of GPA = 0.6

Maximum value of GPA = 4.0

Range = 4.0 - 0.6 = 3.4

Number of classes required = 7

So class width = Range / 7 = 3.4/7 = 0.5

Now sorted data are:

0.6

0.8

1.1

1.2

1.6

1.6

1.6

1.7

1.7

1.9

2.0

2.1

2.1

2.3

2.3

2.4

2.6

2.6

2.6

2.7

2.8

2.9

3.0

3.1

3.2

3.3

3.5

3.5

3.8

3.9

4.0

4.0

Frequency distribution: